3) Assume that X follows a continuous uniform distribution on the interval [3.4, 8.1].

Find E[X].

Provide an exact answer.
Answer #1

Solution :

Given that,

a = 3.4

b = 8.1

Expected value = E(X) = (a + b) / 2

Expected value = E(X) = (3.4 + 8.1) / 2

Expected value = E(X) = 5.75